in a mutual fund with a 7% load and expense ratio of 075%. You can Invest instead in a bank an CD paying 2% interest. a. If you plan to Invest for 4 years, what annual rate of return must the fund portfollo earn for you to be better off in the fund than in the CD? Assume annual compounding of returns. (Do not round Intermediate calculations. Round your answer to 2 declmal places.) b. What annual rate of return must the fund portfolio earn if you plan to invest for 8 years to be better off in the fund than in the CD? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Now suppose that instead of a front-end load the fund assesses a 12b-1 fee of .50% per year what annual rate of return must the fund portfolio earn for you to be better off in the fund than in the CD? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

a. If you invest for 4 years, your CD will give =(1+0.02)^4

If you invest in the mutual fund, you have a 7% loan and a 0.75% expense ratio

(1-0.07)*(1-0.0075)^4 *(1+r)^4 = (1.02)^4

(1+r)^4 = 1.19948

r = 0.0465. 4.65%

Mutual fund should earn 4.66%

b. If you invest for 8 years

(1-0.07)*(1-0.0075)^8*(1+r)^8 = 1.02^8

(1+r)^ 8= 1.3381

r = 0.0371 = 3.71%
Mutual fund should earn 3.71%

c. We cannot answer (c) since we do not know if the investment horizon is 4 years or 8 years. The question is incomplete